OpenStudy (anonymous):

t=p+3r/6. solve for r

OpenStudy (anonymous):

pllzz help me asap. will be much appreciated

OpenStudy (anonymous):

(t-p)/2=r unless those variables stand for something, the equation above is the only solution for r

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

wait how did u solvve it?

OpenStudy (anonymous):

i just rearranged the equation first i subtracted the p and moved it to the left side then i simplified the 3r/6 term to just "r/2" *then i made a mistake :/ it should be (t-p)=r/2 at this point then multiply by 2 to isolate the r **your answer should be 2(t-p)=r